# The tidal evolution of anisotropic subhaloes: A new pathway to creating isotropic and cored satellites

### v1

@@ 3.2 Non-universality of subhalo tidal tracks
Figure 8: Tidal tracks, showing the evolution of the normalised peak circular velocity, $V_{\text{max}}/V_{\text{max,0}}$, against $f_{\rm bound}$ for subhaloes with different initial velocity anisotropy, as indicated in the top-right corner of each panel. Results obtained along different orbits are indicated using different symbols, as indicated in the bottom-right corner of the leftmost panel. Data points of eccentric orbits are computed at apo-center passages. Only isotropic subhaloes follow the well-established, isotropic tidal track (e.g. Green & van den Bosch, 2019) shown as a gray-dashed curve. The tidal tracks of radially (tangentially) biased subhaloes all lie below (above) this isotropic relation. For comparison, the orange-dotted curve shows the tidal track of an isotropic cored system obtained by Peñarrubia et al. (2010).
@@
While tidal tracks may no longer be universal, we now examine whether at least the tidal evolution of subhaloes follows unique tidal tracks for a given anisotropy profile. For that purpose we evolve a setset of our subhaloes (the OM subhalo and those with constant $\beta=0$, $+0.25$, $+0.50$, and $-0.50$) along four different orbits characterised by $(\mathcal{R}_{E},e)=(0.25,0.9),(0.3,0.45),(0.5,0.0),$ and $(0.7,0.9)$. Fig. 8 plots the maximum circular velocity, normalised to that at infall, as a function of the instantaneous bound mass fraction, $f_{\rm bound}$. Results for different orbits are indicates with different symbols, as indicated. For eccentric orbits we follow Peñarrubia et al. (2010), and only show data points close to their apo-centers, where the subhalo is close to quasi-equilibrium. The gray-dashed is the tidal track for an isotropic, NFW subhalo from Green & van den Bosch (2019). As expected, the isotropic subhaloes follow this empirical isotropic tidal track independent of the adopted subhalo orbital parameters, consistent with previous studies (e.g. Hayashi et al., 2003; Peñarrubia et al., 2010; Errani & Navarro, 2021; Amorisco, 2021). The anisotropic subhaloes, on the other hand, clearly deviate from this isotropic tidal track. However, they do follow equally narrow, evolutionary tracks that clearly seem to be independent of the subhalo orbital parameters. Hence, the tidal evolution of subhaloes does trace out tidal tracks, but they depend on the full density and velocity anisotropy profile of the subhalo at infall.
@@
Note how the tidal tracks for the radially and tangentially anisotropic subhaloes fall, respectively, below and above that of the isotropic subhalo. As shown in §4.1, the radially anisotropic subhalo with $\beta=+0.5$ undergoes a rapid, tidal cusp-to-core transformation, and both Peñarrubia et al. (2010) and Errani et al. (2023) have shown that cored subhaloes follow tidal tracks that fall below that of their isotropic counterparts. Interestingly, the tidal track of the $\beta=+0.5$ subhalo (middle panel) closely follows the tidal track for isotropic, cored profiles proposed by Peñarrubia et al. (2010) down to $f_{\rm bound}\sim 0.1$, below which the two evolution tracks diverge. Note also how the tidal track for the OM subhalo initially drops significantly below the isotropic tidal track. However, once the radially biased outskirts have been stripped ($f_{\rm bound}\lesssim 0.05$), leaving a close-to-isotropic cusp, the tidal evolution is ‘back on track’, closely following that of the isotropic subhalo.
@@ 4 Tidal isotropisation and core formation
Fig. 9 compares the time evolution of the density profile (top row), the velocity anisotropy profile (middle), and the 3D velocity dispersion (bottom) for the isotropic subhalo, the radially anisotropic subhaloes with $\beta=+0.25$ and $+0.5$, the tangentially anisotropic subhalo with $\beta=-0.5$, and the OM subhalo, as indicated. In each case, the orbit is the same circular orbit with $\mathcal{R}_{E}=0.25$ as in Figs. 2 and 3 and the top panel of Fig. 4. Different coloured lines correspond to different times, as indicated. For reference, we also mark the tidal radii, $r_{{\rm t}}$, at two of those epochs using vertical lines of matching colour. Here the tidal radius is defined as
@@
Evidently, density profiles exhibit distinctive evolution depending on the corresponding velocity anisotropy at infall. In the case of the tangentially anisotropic subhalo, the central density remains virtually unaffected for the entire duration of the simulation ($25\,{\rm Gyr}$), and the tides only affect (reduce) the density profiles in the outskirts. In the case of both the isotropic and OM subhaloes, the tides cause a slight reduction of the central densities, by about a factor of two. The reduction of the central density is even more pronounced (almost an order of magnitude) for the radially anisotropic subhalo with $\beta=+0.25$, while the $\beta=+0.5$ subhalo reveals a dramatic, and rapid, cusp-to-core transition; already after only $1\,{\rm Gyr}$ (half the subhalo crossing time) the inner density cusp has flattened noticeably, and by $t=6\,{\rm Gyr}$ it has developed a pronounced constant-density core in its center. At this time, the instantaneous tidal radius, (indicated by the light-blue, vertical line) is comparable to the core radius, and the system undergoes complete disruption shortly hereafter.
The panels in the middle row of Fig. 9 show the evolution in the velocity anisotropy profile, $\beta(r)$. The initially isotropic subhalo remains isotropic near the center, but develops a pronounced ‘tangentially biased dip’ around the instantaneous tidal radius, the ‘depth’ of which steadily increases with time. Note how a similar feature is also present in the other subhaloes, with the possible exception of the $\beta=+0.5$ case. The most striking aspect of the evolution of the anisotropy profiles, though, is that in all cases, they become more isotropic over time. In the case of the OM subhalo, the outskirts change from being strongly radially anisotropic to close to isotropic (with the exception of the tangentially anisotropic dip near $r_{{\rm t}}$) in under a Gyr, all the while preserving isotropy in its central region. The two radially anisotropic subhaloes that start out with $\beta=+0.25$ and $+0.50$, respectively, both have basically become completely isotropic, within the noisiness of the measurement and once again with the exception of the tangentially anisotropic dip, after less than $1\,{\rm Gyr}$ of tidal evolution. The tangentially anisotropic subhalo reveals the least amount of evolution, but even that system become close to isotropic in its central region, and again in a very short time. Hence, it appears that being exposed to a strong tidal field causes ‘tidal isotropisation’. We will investigate the cause of this in effect in detail in §4.2 below.
@@
Figure 10: Fraction of constituent particle orbits with peri-centric distance $r_{\text{peri}}<r$ that have an apo-center $r_{\text{apo}}>r_{\text{s,0}}$, colour-coded by initial anisotropy as in Fig. 6. This fraction is indicative of the impact that tidal stripping with $r_{\rm t}=r_{\text{s,0}}$ has on the density at radii $r<r_{\rm t}$. Since more radial orbits have smaller pericenters, stripping of subhaloes that are more radially anisotropic affects the mass distribution down to smaller radii; this is how outside-in stripping of radially anisotropic subhaloes can directly affect the central density, promoting rapid core formation (cf. Figs. 9 and 12).
@@ 4.1 Tidal core formation
One of the key new insights of this study is that the tidal stripping of radially anisotropic subhaloes can cause a rapid cusp-to-core transition. In order to understand the physical origin of this tidal core-formation, it is once again useful to examine the subhalo’s orbital composition. As we have seen in §3.1, since the orbits of more radially anisotropic orbits have larger apo-centers, a larger fraction of the subhalo’s mass will be stripped off for a given tidal radius. Fig. 10 shows the fraction of all particles in the initial conditions with a peri-center $r_{\rm peri}<r$ that have an apo-center $r_{\rm apo}>r_{\text{s,0}}$, plotted as a function of $r$. Envision a tidal field such that the tidal radius $r_{\rm t}=r_{\text{s,0}}$. On a circular orbit this means that, within roughly one dynamical time, all particles with $r_{\text{apo}}>r_{\rm t}=r_{\text{s,0}}$ will be stripped off. Fig. 10 then shows the fraction of those stripped particles that, prior to being stripped, contribute mass down to radii $<r$. Since $r_{\text{peri}}\leq r_{\text{apo}}$, this fraction at $r=r_{\text{s,0}}$ simply reflects the fraction of orbits with $r_{\text{apo}}>r_{\text{s,0}}$, which, as we have already seen in Fig. 6, is larger in systems with larger $\beta$, which explains their higher mass-loss rates. More importantly, as is evident from Fig. 10, systems with larger $\beta$ also contribute more mass at small radii. In particular, in the case with $\beta=+0.5$, roughly 75 percent of all orbits that contribute mass below a scale of $0.01r_{\text{s,0}}$, will be stripped of once the tidal radius reaches down to $r_{\rm t}=r_{\text{s,0}}$. For comparison, in the case with $\beta=0.0$, this fraction is only $\sim 15$ percent, while it is close to zero when $\beta=-0.5$. Hence, stripping of a more radially anisotropic subhalo directly impacts the subhalo further down towards its center. Despite the fact that stripping only operates outside of the tidal radius, it can directly impact (lower) the central density, and the efficiency thereof depends on the peri-centers of the orbits that are being stripped. The further inwards the orbits that are being stripped penetrate, the more strongly the structural integrity of the subhalo’s central density cusp is affected.
@@
Figure 11: Evolution of $\beta_{50\%}$, defined as the velocity anisotropy of the instantaneous $50\%$ most bound particles, for subhaloes colour-coded by their initial velocity anisotropy (as indicated) and evolved on orbits with $(\mathcal{R}_{E},e)=(0.25,0.0)$ (top panel), $(0.25,0.9)$ (middle panel), or in isolation (bottom panel). Note that we have used two different scales for the time-axis in order to better highlight the rapid, leading-order isotropisation ($\beta_{50\%}\rightarrow 0$) within roughly half an orbital period (indicated by the vertical dotted line), and the much slower isotropisation due to tidal shocking that is most pronounced along more eccentric orbits. Note how prior to complete disruption (marked by stars), $\beta_{50\%}(t)$ undergoes large fluctuations. As is evident from the lower panel, systems with large radial anisotropy in their centers undergo a mild level of isotropisation in isolation due to the radial orbit instability. Importantly, this ‘spontaneous’ isotropisation is clearly distinct from, and less pronounced than, the tidal isotropisation evident in the top and middle panels. See text for a detailed discussion.
@@
It is this non-local aspect of tidal stripping, which is more pronounced for more radial orbits, that drives the rapid cusp-to-core transition in systems with sufficiently high radial anisotropy. This is aided by two processes that amplify the impact of the tidal stripping. First of all, as the system re-virialises in response to the mass loss, the system puffs up, which further contributes to lowering the (central) density. In addition, it ‘pushes’ the apo-centers of the remaining particles to larger radii, which increases their probability of being stripped (Kampakoglou & Benson, 2007; Pullen et al., 2014). Secondly, in a central core all particles have similar orbital frequencies, while in a cusp more bound particles typically have higher frequencies. Orbits with higher frequencies are more likely to be adiabatically shielded from impulsive shocks that inject energy into the subhalo during peri-centric passages (Weinberg, 1994; Gnedin et al., 1999). Since cusps are shielded from this energy injection, they typically are immune to tidal shocking, which explains why cusps are so resilient to tidal disruption. Cores on the other hand are typically not shielded, and will experience energy injection which, post re-virialisation, causes the core to expand during each peri-centric passage in the host halo (e.g., Kampakoglou & Benson, 2007; Errani & Peñarrubia, 2020). All these effects conspire to amplify the removal of mass from the central region, ultimately resulting in tidal core-formation. As tidal mass stripping continues, these cored subhalo bound remnants will undergo complete tidal disruption once the instantaneous tidal radius becomes comparable to the core radius (Peñarrubia et al., 2010).
@@
Figure 12: Shell-averaged density profiles of the bound remnants of subhaloes evolved on orbits with $\mathcal{R}_{E}=0.25$ and $e=0.9$ (Fig. 4), colour-coded by the initial anisotopy as in Fig. 1. The leftmost to rightmost columns show time slices at 00 Gyr and the subsequent apo-center passages ($t_{\text{orb}}=1.95$ Gyr), where the bound remnants are close to quasi-equilibrium. Top row: More radially biased subhaloes undergo more rapid and pronounced core formation, due to preferential tidal stripping of constituent radial orbits (Fig. 10). Complete core disruption occurs shortly after the instantaneous tidal radii (vertical lines) become comparable to the respective core radii. The central core density and radius depend on both the initial anisotropy and $f_{\rm bound}$. Middle and Bottom rows: We additionally show profiles in isolation (gray dashed) and at $t=0$ Gyr (black solid). For the isolated subhaloes, the collective-effect-driven isotropisation (cf. Fig. 11) at small radii can reduce the central density slope, but importantly on time and spatial scales distinct from the tide-driven core formation.

### v2

@@ 3.2 Non-universality of subhalo tidal tracks
Figure 8: Tidal tracks, showing the evolution of the normalised peak circular velocity, $V_{\text{max}}/V_{\text{max,0}}$, against $f_{\rm bound}$ for subhaloes with different initial velocity anisotropy, as indicated in the top-right corner of each panel. Results obtained along different orbits are indicated using different symbols, as indicated in the bottom-right corner of the leftmost panel. Data points of eccentric orbits are computed at apo-centre passages. Only isotropic subhaloes follow the well-established, isotropic tidal track (e.g. Green & van den Bosch, 2019) shown as a gray-dashed curve. The tidal tracks of radially (tangentially) biased subhaloes all lie below (above) this isotropic relation. For comparison, the orange-dotted curve shows the tidal track of an isotropic cored system obtained by Peñarrubia et al. (2010).
@@
While tidal tracks may no longer be universal, we now examine whether at least the tidal evolution of subhaloes follows unique tidal tracks for a given anisotropy profile. For that purpose we evolve a setset of our subhaloes (the OM subhalo and those with constant $\beta=0$, +0.25+0.25, +0.50+0.50, and 0.50-0.50) along four different orbits characterised by $(\mathcal{R}_{E},e)=(0.25,0.9),(0.3,0.45),(0.5,0.0),$ and $(0.7,0.9)$. Fig. 8 plots the maximum circular velocity, normalised to that at infall, as a function of the instantaneous bound mass fraction, $f_{\rm bound}$. Results for different orbits are indicates with different symbols, as indicated. For eccentric orbits we follow Peñarrubia et al. (2010), and only show data points close to their apo-centres, where the subhalo is close to quasi-equilibrium. The gray-dashed is the tidal track for an isotropic, NFW subhalo from Green & van den Bosch (2019). As expected, the isotropic subhaloes follow this empirical isotropic tidal track independent of the adopted subhalo orbital parameters, consistent with previous studies (e.g. Hayashi et al., 2003; Peñarrubia et al., 2010; Errani & Navarro, 2021; Amorisco, 2021). The anisotropic subhaloes, on the other hand, clearly deviate from this isotropic tidal track. However, they do follow equally narrow, evolutionary tracks that clearly seem to be independent of the subhalo orbital parameters. Hence, the tidal evolution of subhaloes does trace out tidal tracks, but they depend on the full density and velocity anisotropy profile of the subhalo at infall.
@@
Note how the tidal tracks for the radially and tangentially anisotropic subhaloes fall, respectively, below and above that of the isotropic subhalo. As shown in §4.1, the radially anisotropic subhalo with $\beta=+0.5$ undergoes a rapid, tidal cusp-to-core transformation, and both Peñarrubia et al. (2010) and Errani et al. (2023) have shown that cored subhaloes follow tidal tracks that fall below that of their isotropic counterparts. Interestingly, the tidal track of the $\beta=+0.5$ subhalo (middle panel) closely follows the tidal track for isotropic, cored profiles proposed by Peñarrubia et al. (2010) down to $f_{\rm bound}\sim 0.1$, below which the two evolution tracks diverge. Note also how the tidal track for the OM subhalo initially drops significantly below the isotropic tidal track. However, once the radially biased outskirts have been stripped ($f_{\rm bound}\lesssim 0.12$), leaving a close-to-isotropic cusp, the tidal evolution is ‘back on track’, closely following that of the isotropic subhalo.
@@ 4 Tidal isotropisation and core formation
Fig. 9 compares the time evolution of the density profile (top row), the velocity anisotropy profile (middle), and the 3D velocity dispersion (bottom) for the isotropic subhalo, the radially anisotropic subhaloes with $\beta=+0.25$ and +0.5+0.5, the tangentially anisotropic subhalo with $\beta=-0.5$, and the OM subhalo, as indicated. In each case, the orbit is the same circular orbit with $\mathcal{R}_{E}=0.25$ as in Figs. 2 and 3 and the top panel of Fig. 4. Different coloured lines correspond to different times, as indicated. For reference, we also mark the tidal radii, $r_{{\rm t}}$, at two of those epochs using vertical lines of matching colour. Here the tidal radius is defined as
@@
Evidently, density profiles exhibit distinctive evolution depending on the corresponding velocity anisotropy at infall. In the case of the tangentially anisotropic subhalo, the central density remains virtually unaffected for the entire duration of the simulation ($25\,{\rm Gyr}$), and the tides only affect (reduce) the density profiles in the outskirts. In the case of both the isotropic and OM subhaloes, the tides cause a slight reduction of the central densities, by about a factor of two. The reduction of the central density is even more pronounced (almost an order of magnitude) for the radially anisotropic subhalo with $\beta=+0.25$, while the $\beta=+0.5$ subhalo reveals a dramatic, and rapid, cusp-to-core transition; already after only $1\,{\rm Gyr}$ (half the subhalo crossing time) the inner density cusp has flattened noticeably, and by $t=6\,{\rm Gyr}$ it has developed a pronounced constant-density core in its centre. At this time, the instantaneous tidal radius, (indicated by the light-blue, vertical line) is comparable to the core radius, and the system undergoes complete disruption shortly hereafter.
The panels in the middle row of Fig. 9 show the evolution in the velocity anisotropy profile, $\beta(r)$. The initially isotropic subhalo remains isotropic near the centre, but develops a pronounced ‘tangentially biased dip’ around the instantaneous tidal radius, the ‘depth’ of which first steadily deepens with time and saturates at $\beta\sim-0.3$. Note how a similar feature is also present in the other subhaloes, with the possible exception of the $\beta=+0.5$ case. As shown in Rozier & Errani (2024), such tangential dips are a natural outcome of tide-driven re-virialisation. The most striking aspect of the evolution of the anisotropy profiles, though, is that in all cases, they become more isotropic over time. In the case of the OM subhalo, the outskirts change from being strongly radially anisotropic to close to isotropic (with the exception of the tangentially anisotropic dip near $r_{{\rm t}}$) in under a Gyr, all the while preserving isotropy in its central region. The two radially anisotropic subhaloes that start out with $\beta=+0.25$ and +0.50+0.50, respectively, both have basically become completely isotropic, within the noisiness of the measurement and once again with the exception of the tangentially anisotropic dip, after less than $1\,{\rm Gyr}$ of tidal evolution. The tangentially anisotropic subhalo reveals the least amount of evolution, but even that system become close to isotropic in its central region, and again in a very short time. Hence, it appears that being exposed to a strong tidal field causes ‘tidal isotropisation’. We will investigate the cause of this effect in detail in §4.2 below.
@@
Figure 10: Fraction of constituent particle orbits with peri-centric distance $r_{\text{peri}}<r$ that have an apo-centre $r_{\text{apo}}>r_{\text{s,0}}$, colour-coded by initial anisotropy as in Fig. 6. This fraction is indicative of the impact that tidal stripping with $r_{\rm t}=r_{\text{s,0}}$ has on the density at radii $r<r_{\rm t}$. Since more radial orbits have smaller peri-centres, stripping of subhaloes that are more radially anisotropic affects the mass distribution down to smaller radii; this is how outside-in stripping of radially anisotropic subhaloes can directly affect the central density, promoting rapid core formation (cf. Figs. 9 and 12).
@@ 4.1 Tidal core formation
One of the key new insights of this study is that the tidal stripping of radially anisotropic subhaloes can cause a rapid cusp-to-core transition. In order to understand the physical origin of this tidal core-formation, it is once again useful to examine the subhalo’s orbital composition. As we have seen in §3.1, since the orbits of more radially anisotropic orbits have larger apo-centres, a larger fraction of the subhalo’s mass will be stripped off for a given tidal radius. Fig. 10 shows the fraction of all particles in the initial conditions with a peri-centre $r_{\rm peri}<r$ that have an apo-centre $r_{\rm apo}>r_{\text{s,0}}$, plotted as a function of $r$. Envision a tidal field such that the tidal radius $r_{\rm t}=r_{\text{s,0}}$. On a circular orbit this means that, within roughly one dynamical time, all particles with $r_{\text{apo}}>r_{\rm t}=r_{\text{s,0}}$ will be stripped off. Fig. 10 then shows the fraction of those stripped particles that, prior to being stripped, contribute mass down to radii <r<r. Since $r_{\text{peri}}\leq r_{\text{apo}}$, this fraction at $r=r_{\text{s,0}}$ simply reflects the fraction of orbits with $r_{\text{apo}}>r_{\text{s,0}}$, which, as we have already seen in Fig. 6, is larger in systems with larger $\beta$, which explains their higher mass-loss rates. More importantly, as is evident from Fig. 10, systems with larger $\beta$ also contribute more mass at small radii. In particular, in the case with $\beta=+0.5$, roughly 75 percent of all orbits that contribute mass below a scale of $0.01r_{\text{s,0}}$, will be stripped off once the tidal radius reaches down to $r_{\rm t}=r_{\text{s,0}}$. For comparison, in the case with $\beta=0.0$, this fraction is only $\sim 15$ percent, while it is close to zero when $\beta=-0.5$. Hence, stripping of a more radially anisotropic subhalo directly impacts the subhalo further down towards its centre. Despite the fact that stripping only operates outside of the tidal radius, it can directly impact (lower) the central density, and the efficiency thereof depends on the peri-centres of the orbits that are being stripped. The further inwards the orbits that are being stripped penetrate, the more strongly the structural integrity of the subhalo’s central density cusp is affected.
@@
Figure 11: Evolution of $\beta_{50\%}$, defined as the velocity anisotropy of the instantaneous $50\%$ most bound particles, for subhaloes colour-coded by their initial velocity anisotropy (as indicated) and evolved on orbits with $(\mathcal{R}_{E},e)=(0.25,0.0)$ (top panel), $(0.25,0.9)$ (middle panel), or in isolation (bottom panel). Note that we have used two different scales for the time-axis in order to better highlight the rapid, leading-order isotropisation ($\beta_{50\%}\rightarrow 0$) within roughly half an orbital period (indicated by the vertical dotted line), and the much slower isotropisation due to tidal shocking that is most pronounced along more eccentric orbits. Note how prior to complete disruption (marked by stars), $\beta_{50\%}(t)$ undergoes large fluctuations. As is evident from the lower panel, systems with large radial anisotropy in their centres undergo a mild level of isotropisation in isolation due to the radial orbit instability. Importantly, this ‘spontaneous’ isotropisation is clearly distinct from, and less pronounced than, the tidal isotropisation evident in the top and middle panels. See text for a detailed discussion.
@@
It is this non-local aspect of tidal stripping, which is more pronounced for more radial orbits, that drives the rapid cusp-to-core transition in systems with sufficiently high radial anisotropy. This is aided by two processes that amplify the impact of the tidal stripping. First of all, as the system re-virialises in response to the mass loss, the system puffs up, which further contributes to lowering the (central) density. In addition, it ‘pushes’ the apo-centres of the remaining particles to larger radii, which increases their probability of being stripped (Kampakoglou & Benson, 2007; Pullen et al., 2014). Secondly, in a central core all particles have similar orbital frequencies, while in a cusp more bound particles typically have higher frequencies. Orbits with higher frequencies are more likely to be adiabatically shielded from impulsive shocks that inject energy into the subhalo during peri-centric passages (Weinberg, 1994; Gnedin et al., 1999). Since cusps are shielded from this energy injection, they typically are immune to tidal shocking, which explains why cusps are so resilient to tidal disruption. Cores on the other hand are typically not shielded, and will experience energy injection which, post re-virialisation, causes the core to expand during each peri-centric passage in the host halo (e.g., Kampakoglou & Benson, 2007; Errani & Peñarrubia, 2020). All these effects conspire to amplify the removal of mass from the central region, ultimately resulting in tidal core-formation. As tidal mass stripping continues, these cored subhalo bound remnants will undergo complete tidal disruption once the instantaneous tidal radius becomes comparable to the core radius (Peñarrubia et al., 2010).
@@
Figure 12: Shell-averaged density profiles of the bound remnants of subhaloes evolved on orbits with $\mathcal{R}_{E}=0.25$ and $e=0.9$ (Fig. 4), colour-coded by the initial anisotopy as in Fig. 1. The leftmost to rightmost columns show time slices at 0 Gyr and the subsequent apo-centre passages ($t_{\text{orb}}=1.95$ Gyr), where the bound remnants are close to quasi-equilibrium. Top row: More radially biased subhaloes undergo more rapid and pronounced core formation, due to preferential tidal stripping of constituent radial orbits (Fig. 10). Complete core disruption occurs shortly after the instantaneous tidal radii (vertical lines) become comparable to the respective core radii. The central core density and radius depend on both the initial anisotropy and $f_{\rm bound}$. Middle and Bottom rows: We additionally show profiles in isolation (gray dashed) and at $t=0$ Gyr (black solid). For the isolated subhaloes, the collective-effect-driven isotropisation (cf. Fig. 11) at small radii can reduce the central density slope, but importantly on time and spatial scales distinct from the tide-driven core formation.